A circular loop of radius r is placed in a region where magnetic field increases with respect to time as B(t)=at then induced emf in coil
 

a.  pi r a

b. 3Pi r a

c. 2 pi r 2 a

d. 4 pi r 2 a 

 
 

Grade:12th pass

1 Answers

Arun
4 years ago
Dear student
 
phi = BA
 
E = d(phi)/dt = A dB/dt
 
since B = at, hence dB/dt = a
 
Hence
 
E = a* A = a * pi * r2